OpenStudy (precal):

Does anyone remember how to take the derivative of the absolute value function?

OpenStudy (amistre64):

yep; except for at the tip

OpenStudy (precal):

\[f(x)=1-\left| x \right| \] from interval -1 to 1

OpenStudy (precal):

Isn't it a piecewise function?

OpenStudy (amistre64):

f(x) = x; x>0 = -x; x</0

OpenStudy (amistre64):

piecewise is how im used to it yes

OpenStudy (precal):

Thanks, I thought so. I am doing Rolle's Theorem

OpenStudy (amistre64):

the slope of the abs() depends on which side your on :)

OpenStudy (precal):

I think it fails Rolle's Theorem

OpenStudy (amistre64):

since the tip is a cusp; it might fail it

OpenStudy (precal):

I have to show f(a)=f(b) a=-1 b=1 endpoints of interval

OpenStudy (amistre64):

another prospect is: |x| = sqrt(x^2)

OpenStudy (amistre64):

\[\frac{2x}{2\sqrt{x^2}}\]

OpenStudy (precal):

yes but Rolle's Theorem states f(a)=f(b) on a closed interval [a,b] and differentiable on (a,b), I believe it fails Rolle's Theorem

OpenStudy (amistre64):

we get 0/0 if we try .. so yeah, id agree on intuition

OpenStudy (precal):



OpenStudy (precal):

 fails because you have to be able to differentiate on (a,b)

OpenStudy (amistre64):

if we average the limit from the left and right we get 0 ;)

OpenStudy (precal):

Thanks :) I think I am done with this problem.